### 01 May 2026 v1

# Effect-Transparent Governance for AI Workflow Architectures: Semantic Preservation, Expressive Minimality, and Decidability Boundaries

Abstract
We present a machine-checked formalization of structurally governed AI workflow architectures and prove that effect-level governance can be imposed without reducing internal computational expressivity. Using Interaction Trees in Rocq 8.19, we define a governance operator $\mathcal{G}$ that mediates all effectful directives, including memory access, external calls, and oracle (LLM) queries. Our development compiles with 0 admitted lemmas and consists of 36 modules,  12,000 lines of Rocq, and 454 theorems.
We establish seven properties: (P1) governed Turing completeness, (P2) governed oracle expressivity, (P3) a decidability boundary in which governance predicates are total and closed under Boolean composition while semantic program properties remain non-trivial and undecidable by governance, (P4) goal preservation for permitted executions, (P5) expressive minimality of primitive capabilities (compute, memory, reasoning, external call, observability), (P6) subsumption asymmetry showing structural governance strictly subsumes content-level filtering, and (P7) semantic transparency: on all executions where governance permits, the governed interpretation is observationally equivalent (modulo governance-only events) to the ungoverned interpretation. Together, these results show that governance and computational expressivity are orthogonal dimensions: governance constrains the effect boundary of programs while remaining semantically transparent to internal computation.
1 Introduction
AI workflow systems increasingly act in the world: they query models, store and retrieve state, call external tools, and coordinate distributed services. As these systems become more capable, governance mechanisms are layered onto them to prevent unsafe or unauthorized behavior. A widely held assumption is that governance inevitably trades against capability: restrictions intended to increase safety reduce the space of behaviors the system can exhibit.
This paper formalizes and mechanically verifies an architectural alternative: effect-level structural governance, extending our prior mechanized development McCann (2026d) with semantic transparency, expressive minimality, and decidability boundary results. Instead of modifying internal computation or post-processing outputs, governance is imposed at the effect boundary of the system. Programs are modeled as interaction trees over a directive event type. Governance is a wrapper over event interpretation in the tradition of algebraic effect handlers Plotkin and Pretnar (2009): it mediates each directive with structural checks (e.g. trust and capability predicates) and governance-only instrumentation (e.g. provenance and observability).
The central semantic question is not merely whether governed executions are safe, but whether governance changes what permitted programs compute. Our main contribution is a machine-checked theorem that it does not.
Main Results (informal).
We prove two foundational properties and five supporting results.
  • Safety: governed interpretation enforces effect safety for all programs and handlers.
  • Semantic transparency: for executions in which governance checks succeed, the governed interpretation is observationally equivalent (modulo governance-only events) to the ungoverned interpretation. In particular, permitted programs compute the same partial function and produce the same reachable return values.
We further establish governed expressivity (Turing completeness and an oracle/LLM extension), an explicit decidability boundary and compositional closure for governance predicates, expressive minimality of primitive capabilities, and a strict subsumption separation between structural governance and content-level filtering.
Artifact.
The complete development compiles in Rocq 8.19 with Interaction Trees, comprising 36 modules,  12,000 lines, 454 theorems, and 0 admitted lemmas. The artifact is available at https://github.com/mashin-live/governance-proofs.
Contributions.
  1. 1.
    A machine-checked semantics for effect-governed AI workflows using Interaction Trees.
  2. 2.
    A semantic transparency theorem: governed execution is equivalent to ungoverned execution on all permitted traces (modulo governance-only events).
  3. 3.
    Governed expressive sufficiency: Turing completeness and oracle extension under governance.
  4. 4.
    Expressive minimality: each primitive capability is structurally necessary; removing any primitive strictly reduces expressiveness.
  5. 5.
    A decidability boundary with compositional closure for governance predicates and a separation showing governance cannot decide halting.
  6. 6.
    A strict subsumption theorem: structural governance subsumes content filtering; content filtering does not imply structural safety.
Scope and companion papers.
This paper focuses on expressiveness and semantic transparency results. It builds on the mechanized safety and invariance proofs in McCann (2026d) and extends the structural governance criterion introduced in McCann (2026e). The algebraic semantics paper McCann (2026a) lifts these system-specific results to a parametric framework, showing that any system satisfying three axioms inherits the safety and transparency properties proved here. The certified purity paper McCann (2026b) addresses the practical enforcement of the pure module constraint that the safety theorems assume. The distribution provenance paper McCann (2026c) extends governance to the supply chain with cryptographic provenance.
2 Motivating Example: Governing Effects Without Altering Computation
Consider a workflow that generates a plan via an LLM, stores the plan, invokes an external tool, and returns a result. Schematically:
ITree.bind (trigger (LLMCall prompt)) (fun plan =>
ITree.bind (trigger (MemStore key plan)) (fun _ =>
ITree.bind (trigger (CallMachine api plan)) (fun resp =>
Ret resp)))
This workflow combines internal computation with effectful operations: oracle inference (LLMCall), state (MemStore), and external action (CallMachine). Content-level governance often operates by modifying outputs (filters, rewriting, alignment prompts), which can change the observable behavior of the system even when execution is allowed.
Structural governance instead wraps effect interpretation. For any base handler $h$, the governed handler $\mathcal{G}(h)$ mediates every directive: it checks governance predicates and either delegates to $h$ or denies execution by divergence. This yields two key observations:
  • Effect safety: unauthorized effects cannot occur because all directives are mediated.
  • Conditional semantic preservation: if governance permits the execution, the return value is unchanged (modulo governance-only events).
The remainder of the paper makes these observations precise and proves them for all programs expressible in the architecture.
3 Background
We recall the interaction-tree model and the governance wrapper construction, introduced in McCann (2026d). All definitions are mechanized in Rocq using Interaction Trees Xia et al. (2020); Zakowski et al. (2021).
3.1 Interaction Trees
Programs are coinductive trees over an event type $E$:
$\mathrm{itree}\ E\ R::=\mathrm{Ret}(v)\mid\mathrm{Tau}(t)\mid\mathrm{Vis}(e,k)$
where e:ESe:E\;S and $k:S\to\mathrm{itree}\ E\ R$.
3.2 Directives and Governance Operator
The directive event type $\mathrm{DirectiveE}$ contains constructors for primitive capabilities: computation morphisms, memory operations, oracle calls, external tool calls, and observability events. A handler interprets directives into an effectful semantic domain.
The governance operator $\mathcal{G}$ wraps directive interpretation with pre-checks and post-instrumentation, denying execution on failed checks. This is handler-level effect mediation Plotkin and Pretnar (2009); Bauer and Pretnar (2015): the handler structure is preserved, with governance interposed at interpretation time:
$\mathcal{G}(h)(d)=\text{pre\_checks}\mathbin{>\!\!>\!\!=}\lambda b.\;\text{if }b\text{ then }h(d)\mathbin{>\!\!>\!\!=}\lambda r.\;\text{post\_record}\mathbin{>\!\!>\!\!=}\lambda\_.\;\mathrm{Ret}(r)\text{ else spin}.$
Definition Gov (h : base_handler) : governed_handler :=
fun R (d : DirectiveE R) =>
ITree.bind pre_governance (fun ok =>
if ok then
ITree.bind (lift_io (h R d)) (fun r =>
ITree.bind post_governance (fun _ =>
ret r))
else
ITree.spin).
Here pre_governance emits four $\mathrm{GovE}$ events (trust check, permission check, phase validation, pre-hooks) that each return a boolean, short-circuiting on failure. post_governance emits three $\mathrm{GovE}$ events (guardrails, provenance record, event broadcast). spin is infinite silent divergence. The type $\mathrm{GovIOE}=\mathrm{GovE}+^{\prime}\mathrm{IOE}$ separates governance-only events from real I/O events.
3.3 Safety Predicate
We express safety via a coinductive predicate gov_safe over governed executions, defined using parameterized coinduction Hur et al. (2013). Informally, $\texttt{gov\_safe}(\text{false},t)$ states that $t$ does not perform unauthorized effects; $\texttt{gov\_safe}(\text{true},t)$ tracks the permitted state.
Variant gov_safeF (F : bool -> itree GovIOE R -> Prop)
: bool -> itreeF GovIOE R (itree GovIOE R) -> Prop :=
| GS_Ret : forall allowed r,
gov_safeF F allowed (RetF r)
| GS_Tau : forall allowed t,
F allowed t -> gov_safeF F allowed (TauF t)
| GS_GovE : forall allowed (s : GovernanceStage) k,
(forall b, F true (k b)) ->
gov_safeF F allowed (VisF (inl1 (GovCheck s)) k)
| GS_IOE : forall (e : IOE X) (k : X -> itree GovIOE R),
(forall x, F true (k x)) ->
gov_safeF F true (VisF (inr1 e) k).
Definition gov_safe := paco2 gov_safe_ bot2.
The key asymmetry: GS_GovE applies at any permission level (governance events grant permission), while GS_IOE requires $\text{allowed}=\text{true}$ (I/O requires prior authorization). Bare I/O at false has no matching constructor and is therefore unsafe.
4 Safety and Semantic Transparency
This section presents the two foundational results: universal effect safety and semantic transparency on permitted executions.
4.1 Universal Effect Safety
Theorem 4.1 (Governed Interpretation Safety).
For any handler $h$ and any program $t:\mathrm{itree}\ \mathrm{DirectiveE}\ R$,
$\texttt{gov\_safe}(\text{false},\;\mathrm{interp}(\mathcal{G}(h),t)).$
Theorem 4.1 is uniform over directive types: it applies equally to memory, oracle calls, tool calls, and observability events.
4.2 Semantic Transparency on Permitted Executions
We define a permissiveness condition capturing that governance checks succeed along an execution. Under this condition, governance is observationally transparent modulo governance-only events.
Definition permissive_gov : forall X, GovE X -> X :=
fun X e => match e with GovCheck _ => true end.
Definition permissive_handler (io_h : forall X, IOE X -> X)
: forall X, GovIOE X -> X :=
fun X e =>
match e with
| inl1 ge => permissive_gov X ge
| inr1 ie => io_h X ie
end.
The permissive handler resolves every governance check to true and delegates I/O events to an arbitrary handler io_h. This models the case where governance permits everything. The semantic question: when governance permits, does it alter the result?
Theorem governed_transparency :
forall (h : base_handler) (io_h : forall X, IOE X -> X)
R (t : itree DirectiveE R),
eutt eq
(interp (fun X e => Ret (permissive_handler io_h X e))
(interp (Gov h) t))
(interp (fun X e => Ret (io_h X e))
(interp h t)).
Theorem 4.2 (Semantic Transparency (informal)).
On all executions where governance permits, the governed interpretation is observationally equivalent (modulo governance-only events) to the ungoverned interpretation.
Observational equivalence.
The equivalence relation eutt eq denotes weak bisimulation up to silent transitions ($\mathrm{Tau}$). Governance-only events are interpreted by the permissive handler and do not alter observable I/O behavior. Consequently, semantic transparency establishes equality of the computed partial function on all executions where governance checks succeed. On the left side, the permissive handler interprets governance-only $\mathrm{GovE}$ events to true and erases them; on the right side, no governance events exist. What remains after erasure is identical: the same I/O events, in the same order, producing the same return values.
Corollary governed_same_result :
forall h io_h R (t : itree DirectiveE R) (v : R),
eutt eq (interp (fun X e => Ret (io_h X e))
(interp h t)) (Ret v) ->
eutt eq (interp (fun X e => Ret (permissive_handler io_h X e))
(interp (Gov h) t)) (Ret v).
Corollary governed_goal_preservation :
forall h io_h R (G : R -> Prop) t (v : R),
G v ->
eutt eq (interp (fun X e => Ret (io_h X e))
(interp h t)) (Ret v) ->
eutt eq (interp (fun X e => Ret (permissive_handler io_h X e))
(interp (Gov h) t)) (Ret v).
Denial trades liveness for safety.
Governance preserves safety by converting unauthorized executions into divergence. When a governance check fails, the wrapped directive evaluates to ITree.spin, producing no observable result. The artifact proves that spin is always safe (spin_gov_safe), that binding a continuation to a divergent tree preserves safety (bind_spin_gov_safe), and that spin never equals a return value (spin_not_ret):
Lemma spin_gov_safe :
forall allowed, gov_safe allowed ITree.spin.
Theorem spin_not_ret :
forall R (v : R),
~ eq_itree eq ITree.spin (Ret v).
The system therefore trades liveness (possible termination) for safety (absence of unauthorized effects), but never produces an incorrect return value. Transparency holds only on permitted executions; denial yields divergence rather than semantic equivalence.
5 Governed Expressivity: Turing and Oracle Completeness
We encode register machines Minsky (1967) to establish Turing completeness and extend the encoding with oracle queries to model LLM integration.
5.1 Governed Turing Completeness
Theorem 5.1 (Governed Turing Completeness).
For any handler $h$, register-machine program $p$, and fuel bound,
$\texttt{gov\_safe}(\text{false},\;\mathrm{interp}(\mathcal{G}(h),\;\texttt{translate\_program}(p,\text{fuel},0))).$
Proof.
Directly from Theorem 4.1, since the translation produces an $\mathrm{itree}$ over $\mathrm{DirectiveE}$. ∎
5.2 Governed Oracle Expressivity
We extend the machine with an oracle instruction O_QUERY that emits an LLMCall directive.
Theorem 5.2 (Governed Oracle Expressivity).
For any handler $h$, oracle program $p$, fuel bound, and program counter,
$\texttt{gov\_safe}(\text{false},\;\mathrm{interp}(\mathcal{G}(h),\;\texttt{translate\_oracle\_program}(p,\text{fuel},\text{pc}))).$
Proof.
Again by Theorem 4.1. ∎
6 Decidability Boundary and Compositional Closure
Governance predicates are structural: they are total, syntactic, and closed under composition. In contrast, semantic properties of programs (e.g. halting) remain non-trivial and cannot be decided by governance. This formalizes the boundary concept from McCann (2026e).
6.1 Governance Policies and Closure
We model governance policies as inductive syntax with a total evaluator. We prove closure under Boolean composition (AND/OR/NOT).
Inductive GovPolicy : Type :=
| PolCapability : TrustLevel -> Capability
-> list Capability -> GovPolicy
| PolTrust : TrustLevel -> TrustLevel -> GovPolicy
| PolAnd : GovPolicy -> GovPolicy -> GovPolicy
| PolOr : GovPolicy -> GovPolicy -> GovPolicy
| PolNot : GovPolicy -> GovPolicy.
Fixpoint eval_policy (pol : GovPolicy) : bool :=
match pol with
| PolCapability tl cap dcaps => capability_allowed tl cap dcaps
| PolTrust tl1 tl2 => trust_at_least tl1 tl2
| PolAnd p1 p2 => eval_policy p1 && eval_policy p2
| PolOr p1 p2 => eval_policy p1 $|$ eval_policy p2
| PolNot p => negb (eval_policy p)
end.
Theorem compositional_closure :
(forall p1 p2, eval_policy (PolAnd p1 p2) = true \/
eval_policy (PolAnd p1 p2) = false) /\
(forall p1 p2, eval_policy (PolOr p1 p2) = true \/
eval_policy (PolOr p1 p2) = false) /\
(forall p, eval_policy (PolNot p) = true \/
eval_policy (PolNot p) = false).
6.2 Governance Cannot Decide Halting
We prove a separation: there exist programs that are indistinguishable to governance predicates (same directive class/shape) but differ in halting behavior; thus governance cannot decide halting.
Theorem governance_cannot_decide_halting :
(* Both programs’ first instruction emits MemoryOp *)
(exists params k,
translate_instruction (INC 0 1) =
ITree.bind (trigger (MemoryOp params)) k) /\
(exists params k,
translate_instruction (INC 0 0) =
ITree.bind (trigger (MemoryOp params)) k) /\
(* But one halts and one diverges *)
rm_halts inc_then_halt 2 (0, init_regs) = true /\
(forall n, rm_halts looping_program n (0, init_regs) = false).
Governance predicates operate on the structural properties of individual directives (event type, trust level, capabilities) and not on recursive control flow. Because two programs can present identical directive structure on their first step yet diverge in global behavior, no total governance predicate can decide termination. This is a consequence of Rice’s theorem Rogers (1967) applied to the decidability boundary: governance correctly restricts itself to the structurally decidable side.
Together, these results establish a structural decidability boundary: governance predicates are total and compositional, while semantic program properties remain undecidable.
7 Expressive Minimality of Primitive Capabilities
Surjection of primitives onto a capability taxonomy is not sufficient to rule out redundancy. We therefore prove expressive minimality: each primitive capability is structurally necessary for full expressiveness.
We classify directive events into disjoint capability classes and show strict separations:
  • Memory necessity: register-machine simulation requires memory directives.
  • Reason necessity: oracle extension requires LLMCall.
  • Call necessity: external action morphisms require tool-call directives.
  • Hierarchy: code << memory << reason << call (strict).
Definition is_memory_event {R} (d : DirectiveE R) : bool :=
match d with
| MemoryOp _ | DBOp _ | FileOp _ => true
| _ => false
end.
Definition is_reason_event {R} (d : DirectiveE R) : bool :=
match d with
| LLMCall _ | LLMCallStream _ => true
| _ => false
end.
Definition is_call_event {R} (d : DirectiveE R) : bool :=
match d with
| CallMachine _ | HTTPRequest _ | ExecOp _
| GraphQLRequest _ | WebSocketOp _ | MCPCall _ => true
| _ => false
end.
Lemma event_classes_disjoint : forall R (d : DirectiveE R),
(* each event belongs to exactly one class *)
(is_memory_event d = true /\ is_reason_event d = false
/\ is_call_event d = false /\ is_observe_event d = false)
\/ (* ... three more disjuncts *).
Theorem primitive_minimality :
(* Memory: required for Turing completeness *)
(forall r l, exists params k,
translate_instruction (INC r l) =
ITree.bind (trigger (MemoryOp params)) k) /\
(* Reason: required for oracle extension *)
(forall params, is_reason_event (LLMCall params) = true /\
is_memory_event (LLMCall params) = false) /\
(* Call: required for external invocation *)
(forall params, is_call_event (CallMachine params) = true /\
is_memory_event (CallMachine params) = false /\
is_reason_event (CallMachine params) = false) /\
(* Event classes are pairwise disjoint *)
(forall R (d : DirectiveE R),
(is_memory_event d = true -> is_reason_event d = false) /\
(is_reason_event d = true -> is_memory_event d = false) /\
(is_call_event d = true -> is_memory_event d = false) /\
(* ... all six pairwise implications *) ).
Theorem 7.1 (Strict expressiveness hierarchy).
The primitive capabilities form a strict chain:
$\textsc{code}\;<\;\textsc{code}+\textsc{memory}\;<\;\textsc{code}+\textsc{memory}+\textsc{reason}\;<\;\textsc{code}+\textsc{memory}+\textsc{reason}+\textsc{call}$
where each level strictly extends the previous: the new primitive emits events that no combination of the existing primitives can produce.
Proof.
Each strict separation combines two facts: (1) the translation at the new level structurally emits events from the new class (e.g. INC emits MemoryOp, O_QUERY emits LLMCall), and (2) event class disjointness ensures the lower level cannot produce those events. ∎
Remark 7.2 (Expressive minimality).
Removing any primitive capability strictly reduces expressive power. This follows from the hierarchy and disjointness results above.
8 Subsumption Asymmetry: Structural vs. Content Governance
Structural governance mediates effects at the interpretation boundary, whereas content governance modifies returned values. The safety predicate $\texttt{gov\_safe}(\text{allowed},t)$ is defined (via paco) as the greatest fixed point of gov_safeF, and the only constructor that admits visible I/O is:
| GS_IOE : (forall x, F true (k x)) ->
gov_safeF F true (VisF (inr1 e) k)
In particular, GS_IOE requires $\text{allowed}=\text{true}$. There is no constructor of gov_safeF that can derive $\texttt{gov\_safeF}\;F\;\text{false}\;(\text{VisF}\;(\text{inr1}\;e)\;k)$.
Therefore, any tree whose current observation is a bare I/O event at permission level false has no derivation of gov_safe. One proves $\lnot\;\texttt{gov\_safe}(\text{false},\;\mathrm{Vis}(\text{inr1}\;e,k))$ by unfolding gov_safe once (via punfold) and inverting on the resulting gov_safeF goal: the $\text{VisF}\;(\text{inr1}\;e)\;k$ case can only be discharged when $\text{allowed}=\text{true}$.
This argument is independent of any content-level filtering in the underlying handler: content governance can change values but does not prevent the occurrence of an unauthorized I/O step. Hence structural governance strictly subsumes content governance: $\mathcal{G}$ enforces mediation of all effectful directives by construction, while content filters alone cannot imply effect safety.
Theorem 8.1 (Structural subsumes content).
For any handler $h^{\prime}$ (including content-filtering handlers) and any program $t$,
$\texttt{gov\_safe}(\text{false},\;\mathrm{interp}(\mathcal{G}(h^{\prime}),t)).$
Theorem 8.2 (Content does not subsume structural).
There exist effectful trees that are content-filtered but not structurally safe.
Corollary 8.3 (Subsumption asymmetry).
Structural governance strictly subsumes content governance.
9 Capstone Theorem
We assemble the properties into a single theorem governed_cognitive_completeness.
Theorem governed_cognitive_completeness :
forall (h : base_handler),
(* P1: Turing complete AND governed *)
(forall (p : program) (fuel : nat),
gov_safe false
(interp (Gov h) (translate_program p fuel 0)))
(* P2: Oracle complete AND governed *)
(forall (p : oracle_program) (fuel : nat) (pc : label),
gov_safe false
(interp (Gov h) (translate_oracle_program p fuel pc)))
(* P3: structural_decidable /\ compositional_closure
/\ halting_nontrivial
/\ governance_cannot_decide_halting *)
... (* 10 conjuncts; see Rice.v *)
(* P4: governed_goal_preservation *)
... (* see GoalDirected.v *)
(* P5: cognitive_surjection /\ primitive_minimality
/\ strict_hierarchy *)
... (* see CognitiveArchitecture.v,
ExpressiveMinimality.v *)
(* P6: structural_subsumes_content
/\ content_not_subsumes_structural *)
... (* see Subsumption.v *)
(* P7: Semantic transparency *)
(forall (io_h : forall X, IOE X -> X) R
(t : itree DirectiveE R),
eutt eq
(interp (fun X e => Ret (permissive_handler io_h X e))
(interp (Gov h) t))
(interp (fun X e => Ret (io_h X e))
(interp h t))).
Structure.
The conjuncts correspond to: governed expressivity (P1–P2), the decidability boundary and closure (P3), goal preservation (P4), expressive minimality (P5), subsumption asymmetry (P6), and semantic transparency (P7). Each conjunct is proved by importing the relevant module theorem and applying it directly.
10 Discussion and Trust Boundary
Our theorems establish properties of a formal model: interaction-tree programs, a governance wrapper, and a safety predicate. The guarantees hold relative to a standard trusted computing base: the Rocq kernel (type-checking the proofs), the Interaction Trees library (providing the coinductive framework), and any runtime that faithfully implements the $\mathcal{G}$ operator and directive interpretation. This TCB is comparable to that of verified compilers Leroy (2009) and verified OS kernels Klein et al. (2009): formal properties hold of the model, and the deployment gap is bridged by testing. In our case, the connection from model to a deployed runtime is validated via a verified interpreter specification tested against over 70,000 randomly generated directive sequences with zero disagreements (included in the artifact). Effect separation for the executor layer is developed in McCann (2026b).
The key design principle established here is not “governance by output restriction,” but effect-transparent governance: constrain the effect boundary while preserving semantics on permitted executions.
Limitations.
Semantic transparency is proved relative to executions in which governance checks succeed; when a check fails, the governed semantics diverges (spin) rather than simulating the ungoverned run. The expressivity results use a fuel-indexed register-machine translation, which captures partiality but abstracts from resource exhaustion, scheduling, and real-time constraints. Oracle calls are modeled as directives with unconstrained responses; we do not formalize probabilistic model behavior, distributional assumptions, or quantitative uncertainty. Governance predicates are intentionally structural and total; the development does not attempt to decide semantic properties such as termination or functional correctness. The trusted computing base includes the Rocq kernel, the Interaction Trees library, and any extraction or interpretation layer used to realize the semantics in a deployed runtime. Accordingly, the theorems are statements about the abstract architecture and its interpreter, not end-to-end guarantees about compiled binaries. Finally, transparency is stated modulo erasure of governance-only events and weak bisimulation (eutt), reflecting preservation of observable results rather than literal trace identity.
11 Related Work
Algebraic effect handlers.
Plotkin and Pretnar Plotkin and Pretnar (2009) introduced handlers for algebraic effects; Bauer and Pretnar Bauer and Pretnar (2015) developed the programming model. Our governance operator is a handler wrapper in the same tradition: it interposes on effect interpretation without modifying the underlying handler. Wadler’s monadic approach Wadler (1995) gives the semantic foundation for interpreting effectful programs through bind; our interaction-tree development is inherently monadic.
Interaction Trees and coinduction.
Xia et al. Xia et al. (2020) introduced interaction trees for representing recursive and impure programs in Rocq; Zakowski et al. Zakowski et al. (2021) applied them to verified LLVM semantics. Hur et al. Hur et al. (2013) developed parameterized coinduction (paco), which underlies both our safety and transparency proofs.
Verified systems.
CompCert Leroy (2009) and seL4 Klein et al. (2009) demonstrate that large formal verification efforts yield reliable systems. Our transparency theorem is structurally analogous to CompCert’s forward simulation: both show that a transformation preserves observable behavior.
Content governance and alignment.
RLHF Ouyang et al. (2022) and Constitutional AI Bai et al. (2022) operate at the output level. Our subsumption theorem formalizes that such methods do not imply structural effect safety.
Verified AI and agent safety.
Prior frameworks propose properties to verify Seshia et al. (2016); Dalrymple and others (2024); we deliver a mechanized development with semantic transparency and expressive minimality.
Own prior work.
The governance model, safety predicate, and Turing completeness proof are given in McCann (2026d). The structural argument for why behavioral governance fails at the effect level is developed in McCann (2026e). Effect separation for the executor layer is formalized in McCann (2026b). The present paper extends these with semantic transparency, expressive minimality, and a decidability boundary with compositional closure.
Cognitive architectures.
Park et al. Park et al. (2023) demonstrated generative agents with perceive-reflect-act loops; Sumers et al. Sumers et al. (2024) surveyed cognitive architectures for language agents. Our primitive capability taxonomy (compute, memory, reason, call, observe) is a formalization of capabilities these architectures require, with machine-checked minimality.
12 Conclusion
We have presented a machine-checked semantics for effect-governed AI workflow architectures and proved that structural governance can be imposed without reducing computational expressivity. Our main theorem shows semantic transparency on permitted executions: governed and ungoverned interpretations are observationally equivalent modulo governance-only events. Supporting results establish governed expressivity (Turing and oracle), a compositional decidability boundary, expressive minimality of primitive capabilities, and strict subsumption of content filtering by structural governance.
Artifact availability.
The complete Rocq development (36 modules,  12,000 lines, 454 theorems, 0 admitted lemmas) is available at https://github.com/mashin-live/governance-proofs.